OpenStudy (anonymous):

a ≠b and 1/x+ 1/a = 1/b then x= A. 1/b= 1/a B. b-a C. 1/ab D. a-b/ab E. ab/a-b

OpenStudy (anonymous):

\[\frac{ 1 }{ x }=\frac{ 1 }{ b }-\frac{ 1 }{ a }=\frac{ a-b }{ ab },x=?\]

OpenStudy (anonymous):

Where to start reading it from?

OpenStudy (anonymous):

\[\frac{1}{x}+\frac{1}{a}=\frac{1}{b}\]?

OpenStudy (anonymous):

subtract \(\frac{1}{a}\) from both sides as @surjithaer wrote

OpenStudy (anonymous):

you get \[\frac{1}{x}=\frac{1}{b}-\frac{1}{a}\] then do the subtraction do you know how to subtract \[\frac{1}{a}-\frac{1}{b}\]?

OpenStudy (anonymous):

is either E or D then?

OpenStudy (anonymous):

E or D are letters you have to subtract

OpenStudy (anonymous):

then is A

OpenStudy (anonymous):

thanks <3

OpenStudy (anonymous):

A doesn't even make sense you are solving for \(x\) and A says \[\frac{1}{b}=\frac{1}{a}\]

OpenStudy (anonymous):

it has to be c then